OpenStudy (anonymous):

solve the equation 4y+5=-31. check your answer.

OpenStudy (anonymous):

so 13 right

OpenStudy (anonymous):

that's what I got

OpenStudy (anonymous):

13

OpenStudy (lazyboy):

wait

OpenStudy (anonymous):

ok

OpenStudy (lazyboy):

Solution: 4y + 5 = ? 31 4y = -31 - 5 4y = -36 y = -36/4 y = -9 (answer) Checking : by substituting value of y 4y + 5 = -31 4 (-9) + 5 = -31 -36 + 5 = - 31 -31 = -31

OpenStudy (lazyboy):

does this help